Circular No. 7665

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2001cy IN UGC 11927 R. Chornock, M. Modjaz, and A. V. Filippenko, University of Cali-

fornia, Berkeley, write: “Inspection of CCD spectra (range 330–1000 nm) taken on July 16 with the Shane 3-m reflector at Lick Observatory shows that SN 2001cy (IAUC 7655) is a type-II supernova soon after maximum light. The spectrum exhibits well-developed H Balmer and Fe II lines. The Hα profile is unusual, with two absorption minima in addition to a broad emission component, similar to the peculiar SN 1999gk (IAUC 7332).”

V1494 AQUILAE M. Bos, Auckland, New Zealand; A. Retter, Keele University; L. Cook,

Concord, CA; and R. Novák, Copernicus Observatory, Brno, report: “Unfiltered and R-band CCD photometry of V1494 Aql (N Aql 1999 No. 2) obtained on 12 nights during June–July shows a robust change in the shape and amplitude of the 0.13467-d period (IAUC 7537). It now has an eclipse shape with depth ∼ 0.5 mag (10 times larger than in mid-2000). A second shallow eclipse (∼ 0.1 mag deep) at phase 0.5 can also be seen. The light curve may be explained by a two-component model. The non-eclipsing source, presumably the white dwarf, has faded significantly, while the luminosity of the eclipsing component (probably the accretion disk, obscured by the companion) nearly stayed constant during the observations. This interpretation indicates a partial obscuration of the disk, and it is thus predicted that the eclipse depth will not exceed 0.75 mag (50 percent) with a further fading of the nova.”

BL LACERTAE J. A. Mattei, AAVSO, reports that this blazar prototype, which bright-

ens almost every year, has brightened again. Observations in the AAVSO International Database indicate that the most recent brightenings have been in July 1997, May 1998, Oct. 1999, and Oct. 2000. The rise to maximum may take several months, and maximum (at average visual mag 13.7 and showing fluctuations) may last 2–3 months. At minimum, BL Lac is between visual mag 15.0 and 16.0. Recent AAVSO magnitudes: Apr. 21.13 UT, 14.7 (G. Poyner, West Midlands, England); May 30.03, 14.9 (Poyner); June 8.33, 13.7 (M. Simonsen, Macomb, MI); 15.95, 13.8 (P. Dubovsky, Podbiel, Slovakia); 29.32, 13.8 (Simonsen); July 5.02, 14.4 (Poyner); 9.28, 13.3 (Simonsen); 10.92, 13.0 (Dubovsky); 12.35, 13.4 (Simonsen); 14.97, 13.9 (Dubovsky); 15.34, 13.9 (Simonsen).

2001 July 17  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2001 CBAT  Daniel W. E. Green